Previous

Calendar

September

2008

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

31

JD244

1

JD245
CU holiday (Labor Day)

2

JD246
Site Visit (JB, NT): Details , Photos (6)

N2 @ 300 psi, span @ 1700 psi

swapped 15 flasks at the west robot

3

JD247

4

JD248
JB at site (Merced day)

span @ 1700 psi

swapped N2 tanks at the tower. retired N2 cc193097 and started cc1255575 @ 1925 psi

SpB Returns

5

JD249

6

JD250

7

JD251

8

JD252
Upgrade OS on Porter (contact= Erinn)

9

JD253
Site Visit (JB, NT): Details

swapped 15 flasks at west robot

downloaded from uc, ucb, cnr, prof, met, nctc, soil, tcs

re-deployed the ATI 2-D sonic anemometer

span @ 1690 psi, N2 @ 1800 psi

hauled an old span cylinder and an old N2 cylinder back to the trailer, hauled a new span cylinder out to the tower

10

JD254

11

JD255
SpB in Utah

12

JD256
SpB in Utah

13

JD257

14

JD258

15

JD259

16

JD260
Site Visit (JB, SpB, NT): Details , Photos (23)

hauled last N2 cylinder from the trailer out to the tower

from XX:XX-XX:XX MST, Changed profiler (li-6251) dessicant

tested out calibration of LI-7500 (using LI-610)

re-made TC connected to soil logger (am25t mux, chan_20)

swapped 15 flasks at the west robot

17

JD261
Site Visit (SO): Details

Steve Oncley made some small adjustments to the alignment of the mini-tower (this might have been done a week earlier).

18

JD262

19

JD263

20

JD264

21

JD265

22

JD266
Autumnal equinox

23

JD267
Site Visit (JB, SpB, NT): Details , Photos (24)

from 10:10-11:50 MST, Replaced tubing, heat exchanger and Nu-Pro filters in the LI-6262 flux system

Tested a new snow temperature probe to the cr10 logger (note the wiring needed re-doing)

swapped 15 flasks at the west robot

Clear and cool (classic fall day). The Aspens around the site are at near peak color.

24

JD268

25

JD269

26

JD270

27

JD271
Chan 204 (MET logger) losing 1-hz data (SC932 seems to be up and down)

28

JD272
Chan 204 (MET logger) losing 1-hz data (up and down)

29

JD273
Chan 204 (MET logger) losing 1-hz data (up and down)

30

JD274
Site Visit (JB, NT, SpB): Details , Photos (28)

span @ 1450, N2 @ 1200

swapped 15 flasks at west robot

added 2 accelerometers to the "fast" logger

started running ops47 (added 2 accelerometers)

started digging holes for the CS616 probes

1

JD275
Site Visit (JB, NT, SpB): Details , Photos (12)

tested the replacement laptop at the tower

tested some CS616 probes with the soil logger (see details)

2

JD276

3

JD277
Site Visit (JB, SpB, MP, NT, DB, Diego, SO): Details , Photos (11)

JB, SpB and MP finished removing the PAR wands from grove 1 and started work on grove 2, hauled the adam and some other par equipment back to the trailer

SpB troubleshot the CS616s some more...

4

JD278